40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Aug 12 18:32:37 2010, Alastair, Laser, GYRO, Beam overlap etc TEK00000.PNGTEK00001.PNGTEK00002.PNG
    Reply  Fri Aug 13 15:50:39 2010, Jenna, Laser, GYRO, Beam overlap etc 
Message ID: 932     Entry time: Fri Aug 13 15:50:39 2010     In reply to: 927
Author: Jenna 
Type: Laser 
Category: GYRO 
Subject: Beam overlap etc 


The error signal from the CW side looked much improved after changing the polarization (TEK00001.PNG) but the CW mode still looked unstable.  When we lock both directions of the cavity and look at these signals again we see this (TEK00002.PNG) oscillation in the signal.  It is massive.  It basically takes up the full error signal and is no doubt the reason why the mode looks unstable.

If you turn off the FM input on the VCO driving the AOM then the oscillation is 50% smalller but still there.  We wondered if this was coming from the Tektronix, so we swapped the AOM over to the Marconi, but we see exactly the same noise with this if the FM deviation is set to ~300kHz.  It does go lower if we lower the deviation setting on the Marconi below about 100kHz.  This might be the noise we're seeing in the 95MHz signal in transmission.

**Edit** we just spoke to Frank who thinks this may be because we have the gain too high in this loop (though turning down the PDH box to almost zero didn't get rid of the oscillation).  Tomorrow we'll try reducing this further.

 We were investigating the source of the oscillation in the CW signal when we found that the CCW photodiode was saturating, which must have happened when we were adjusting the power between the two beams. We put in a beamsplitter to reduce the power going onto the PD, and now the CW error signal is no longer oscillating so that seems to have been the source of the problem.

Our 95MHz signal is now at -28dBV, which is 1.3uW of laser power out of 42.5uW (3% contrast ratio), and the DC signal from the photodiode is 30mV which corresponds to 40uW of power on the PD out of 42.5uW, so we're getting most of the power onto the PD now.

ELOG V3.1.3-